Find f(t). 3s F-{ (s - 732 f(t) = 3e7t + 9tet X
Answer #1

To find f(t)

Take partial fraction and do inverse Laplace transform.

Formula used:

L^{-1}[\frac{1}{(s-a)}] =e^{at}

L^{-1}[\frac{1}{(s-a)^2}] =t.e^{at}

Find f(t) 35 1[13 (S-1) 3) 48-49,] (S-4)? Take partial Freation on S (S-1) ? S А B 3 + (S-1) s-1 put S = A(S-1) +Bin S-0 OF31 +31 FIH) +21540 Now take innvrne laplace transform, 31 Cs (5-y)? -Est 16 f(t):341t + 7x3 tett. 30th fit) = 274 ( 3421t).